OpenStudy (anonymous):

Change 4x - 3y = -15 to slope-intercept form and state the slope and y-intercept.

OpenStudy (anonymous):

slope-intercept form: y=4/3X +15/4 slope: m=4/3

OpenStudy (anonymous):

slope intercept form is : y = mx+c so whenever u come across such question convert the equation in this form then you will be able to find the slope

OpenStudy (anonymous):

yeah, you're right. .